OpenStudy (anonymous):

show how to solve 7/9 times 5/14 with and without cancellation

OpenStudy (lgbasallote):

\[\frac 79 \times \frac 5{14} \implies \frac{7 \times 5}{9 \times 14}\] does that help?

OpenStudy (anonymous):

i'm still a bit confused

OpenStudy (lgbasallote):

with what? confused about what i did? or confused about what to do next?